# Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ and Noether-Lefschetz theory of $\mathcal{A}_{g}$

### v1

@@ Abstract.
We calculate the genus 1 Gromov-Witten theory of the Hilbert scheme $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ of points in the plane. The fundamental 1-point invariant (with a divisor insertion) is calculated using a correspondence with the families local curve Gromov-Witten theory over the moduli space ¯$\overline{\mathcal{M}}_{1,1}$. The answer exactly matches a parallel calculation related to the Noether-Lefschetz geometry of the moduli space $\mathcal{A}_{g}$ of principally polarized abelian varieties. As a consequence, we prove that the associated cycle classes satisfy a homomorphism property for the projection operator on $\mathsf{CH}^{*}(\mathcal{A}_{g})$. The fundamental 1-point invariant determines the full genus 1 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ modulo a nondegeneracy conjecture about the quantum cohomology. A table of calculations is given.
@@ 0.1. Hilbert schemes
An open dense set of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ parameterizes ideals associated to configurations of $n$ distinct unordered points. The geometry of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ has been studied from many points of view for several decades now, see [11, 16, 17, 20, 32]. Our perspective here is related to the interactions of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ with Gromov-Witten theory and the relative Donaldson-Thomas invariants of threefolds as developed in [3, 27, 28, 29, 30, 37, 38, 39, 43, 44, 45].
@@
where $[V_{\mu}]$ is (the cohomological dual of) the class of the subvariety of $\mathsf{Hilb}^{|\mu|}(\mathbb{C}^{2})$ with generic element given by a union of schemes of lengths
@@
with fiber $\mathbb{C}[x,y]/\mathcal{I}$ over $[\mathcal{I}]\in\mathsf{Hilb}^{n}(\mathbb{C}^{2})$. The $\mathsf{T}$-action on $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ lifts canonically to the tautological bundle (0.1). Let
@@
The operator $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}$ is calculated explicitly in the Nakajima basis for all $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ in [37]. The matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}$ lie in the field of rational functions in $q$ (with coefficients [^6] In the context of quantum cohomology, the definitions require localization, so we will always consider $H^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ and $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ as modules over $\mathbb{Q}(t_{1},t_{2})$. in $\mathbb{Q}(t_{1},t_{2})$) and determine all genus 00 Gromov-Witten invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ by [37, Section 4.2].
@@
The series (0.2) is always a rational function in $q$,
$\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\in\mathbb{Q}(t_{1},t_{2})(q).\,$
@@ 0.2. Gromov-Witten theory of families of local elliptic curves
Let ¯$\overline{\mathcal{M}}_{1,r}$ be the moduli space of Deligne-Mumford stable curves of genus $1$ with $r$ markings. [^7] We will always assume $r>0$ for stability. Let
@@
be the universal local curve over ¯$\overline{\mathcal{M}}_{1,r}$. The torus $\mathsf{T}=(\mathbb{C}^{*})^{2}$ acts on the $\mathbb{C}^{2}$ factor as before.
Let $\mu^{1},\ldots,\mu^{r}\in\mathrm{Part}(n)$, and let ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$ be the moduli space of stable [^8] The superscript $\bullet$ indicates possibly disconnected domain curves (but no connected component of the domain is contracted to a point). We follow the conventions of [3]. relative maps to the fibers of $\pi_{\mathbb{C}^{2}}$,
@@
is the moduli space of stable maps of genus $g$ to $E\times\mathbb{C}^{2}$ relative to the divisors determined by the nodes and the markings of $E$ with boundary [^9] The boundary conditions are unordered, and the cohomology weights of the boundary conditions are all the identity class. condition $\mu^{i}$ over the divisor $p_{i}\times\mathbb{C}^{2}$. Since the degree $n$ is recorded in the size of the partitions $\mu^{i}$, we omit $n$ from the notation for ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$.
The moduli space ¯$\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$ has $\pi_{\mathbb{C}^{2}}$-relative virtual dimension
$-nr+\sum_{i=1}^{r}\ell(\mu^{i})\,.$
@@
The moduli space of stable maps ¯$\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}$ is empty unless $g[b]$ is an integer. The virtual class [¯$\Big{[}\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}$ is the $\pi_{\mathbb{C}^{2}}$-relative $\mathsf{T}$-equivariant virtual class of the family of relative stable maps to the fibers of $\pi_{\mathbb{C}^{2}}$. We define
@@
$=$ ¯$\int_{\overline{\mathcal{M}}_{1,r}}\epsilon_{*}\left(\Big{[}\overline{\mathcal{M}}^{\bullet}_{g}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}\right).$
@@
after the variable change $-q=e^{iu}$.
@@ 0.3. Moduli of abelian varieties
be the universal principally polarized abelian variety. We refer the reader to [2] for the foundations of the study of the moduli of abelian varieties.
A general abelian variety $(X,\theta)$ parameterized by $\mathcal{A}_{g}$ has Picard number $1$. The Noether-Lefschetz locus of $\mathcal{A}_{g}$ parameterizes abelian varieties with Picard number at least 2. The simplest components of the Noether-Lefschetz locus of $\mathcal{A}_{g}$ are related to the geometry of elliptic curves on abelian varieties. For $g\geq 1$ and $n\geq 1$, let
$\mathsf{NL}_{g,n}=\left\{(X,\theta)\in\mathcal{A}_{g}\left|\begin{array}[]{c}X\text{ contains a subgroup }E\subset X\\ \text{ which is an elliptic curve of degree }\theta\cdot[E]=n\end{array}\right.\right\}.$
Let $[{\mathsf{NL}}_{g,n}]\in\mathsf{CH}^{g-1}(\mathcal{A}_{g})$ be the associated cycle class.101010 The $g=1$ case is degenerate: $[\mathsf{NL}_{1,n}]=[\mathcal{A}_{1}]$ by definition. All Chow groups are taken with $\mathbb{Q}$-coefficients. The following linear combination of components plays a geometrically important role [^11] As usual, $\sigma_{r}(n)=\sum_{d|n}d^{r}$.
$[\widetilde{\mathsf{NL}}_{g,n}]=\sum_{n^{\prime}|n}\sigma_{1}\left(\frac{n}{n^{\prime}}\right)[\mathsf{NL}_{g,n^{\prime}}]\ \in\ \mathsf{CH}^{g-1}(\mathcal{A}_{g})\,,$
@@
or, equivalently [^12] Equation (0.4) is correct also for $g=1$ with the convention $[\mathsf{NL}_{1,n}]=[\mathcal{A}_{1}]$.,
@@ 0.4.1. Hilbert schemes of points
$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(-5-7q-q^{2}+2q^{3}-q^{4}-7q^{5}-10q^{6}-7q^{7}+...\right),$
@@
$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(-\frac{35}{2}-21q-q^{2}-3q^{3}-17q^{4}-21q^{5}-19q^{6}-21q^{7}+...\right)\,,$
@@
$=$ $-\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\left(\frac{136}{3}+\frac{277}{6}q-\frac{41}{6}q^{2}+\frac{17}{3}q^{3}+\frac{151}{6}q^{4}+\frac{127}{6}q^{5}+\frac{101}{3}q^{6}+\frac{277}{6}q^{7}+...\right).$
@@ 0.4.2. Families of local elliptic curves
Let ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)$ be the moduli space of stable [^13] The superscript $\circ$ indicates connected domain curves. relative maps to the fibers of $\pi$,
@@
is the moduli space of stable maps of 1-pointed connected genus $g$ curves to $E$ of degree $n$ relative to the divisors determined by the nodes of $C$. The moduli space ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)$ has $\pi$-relative virtual dimension $12g-$ 1 and total virtual dimension $2g$.
@@ 0.4.3. The homomorphism question
Let $\mathcal{M}_{g}^{\mathsf{ct}}$ be the $33g-$ 3 dimensional moduli space of genus $g$ curves of compact type, and let
@@ 0.5. Reduction of multi-point series to 1-point series
By [37, Section 4.2], the quantum powers of $D$ generate the quantum cohomology $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$. More precisely, the set
@@
is a basis of $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ as a $\mathbb{Q}(t_{1},t_{2})(q)$-vector space. Our first reduction result is that the genus 1 Gromov-Witten theory (0.2) of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ can be effectively reduced to 1-point series in the basis (0.7).
@@ 0.6. Reduction to $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$
The reduction of the genus 1 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to the basic series $\langle D\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ is more subtle. Since the quantum cohomology of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is semisimple, the Givental-Teleman reconstruction result [15, 47] reduces the full higher genus Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to the genus 0 theory together with the calculation of the $\mathsf{R}$-matrix. The $\mathsf{R}$-matrix was proven to be determined [^16] Since the Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is equivariant, the associated $\mathsf{CohFT}$ is not conformal. by constant map invariants in [44]. The resulting control of the full Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ was used in [44] to prove both the crepant resolution and the GW/DT correspondences associated to the geometry, but was not sufficient to obtain closed form evaluations of any higher genus series (because of the difficulty in the determination and the lack of closed forms for the $\mathsf{R}$-matrix).

### v2

@@ Abstract.
We calculate the genus 1 Gromov-Witten theory of the Hilbert scheme $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ of points in the plane. The fundamental 1-point invariant (with a divisor insertion) is calculated using a correspondence with the families local curve Gromov-Witten theory over the moduli space $\overline{\mathcal{M}}_{1,1}$. The answer exactly matches a parallel calculation related to the Noether-Lefschetz geometry of the moduli space $\mathcal{A}_{g}$ of principally polarized abelian varieties. As a consequence, we prove that the associated cycle classes satisfy a homomorphism property for the projection operator on $\mathsf{CH}^{*}(\mathcal{A}_{g})$. The fundamental 1-point invariant determines the full genus 1 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ modulo a nondegeneracy conjecture about the quantum cohomology. A table of calculations is given.
@@ 0.1. Hilbert schemes
An open dense subset of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ parameterizes ideals associated to configurations of $n$ distinct unordered points. The geometry of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ has been studied from many points of view for several decades now, see [12, 17, 18, 21, 34]. Our perspective here is related to the interactions of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ with Gromov-Witten theory and the relative Donaldson-Thomas invariants of threefolds as developed in [4, 28, 29, 31, 32, 39, 40, 41, 45, 46, 47].
@@
where $V_{\mu}]$ is (the cohomological dual of) the class of the subvariety of $\mathsf{Hilb}^{|\mu|}(\mathbb{C}^{2})$ with generic element given by a union of schemes of lengths
@@
with fiber $\mathbb{C}[x,y]/\mathcal{I}$ over $\mathcal{I}]\in\mathsf{Hilb}^{n}(\mathbb{C}^{2})$. The $\mathsf{T}$-action on $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ lifts canonically to the tautological bundle (0.1). Let
@@ 0.2. Quantum cohomology
The operator $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}$ is calculated explicitly in the Nakajima basis for all $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ in [39]. The matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}$ lie in the field of rational functions in $q$ (with coefficients [^6] In the context of quantum cohomology, the definitions require localization, so we will always consider $H^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ and $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ as modules over $\mathbb{Q}(t_{1},t_{2})$. in $\mathbb{Q}(t_{1},t_{2})$) and determine all genus 0 Gromov-Witten invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ by [39, Section 4.2].
@@
The series (0.2) is always a rational [^7] There is a gap in the proof of rationality in the published paper [46] which is filled in a revision available on the authors’ websites. The gap does not affect the proof in [46] of the triangle of correspondences [46, Theorem 2] in the form of analytic continuation from $q=0$ to q=1q=-1, which is all that is needed here. function [46] in $q$,
$\big{\langle}\mu^{1},\mu^{2},\ldots,\mu^{r}\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\in\mathbb{Q}(t_{1},t_{2},q)\,.\,$
@@
The genus 0 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is a very rich subject with connections to the quantum Calogero-Sutherland many body system [39], to the crepant resolution conjecture [3], and to quantum groups [30]. The question of computing $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ for all $n$ was posed in [44, 46]. We will see that the genus 1 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ provides new connections: the theory is fundamentally linked to the geometry of the moduli space of abelian varieties.
There are several natural open questions. What is the structure of the genus 1 Gromov-Witten theory of more general Nakajima quiver varieties? Is there an exact calculation of the full genus Gromov-Witten theory?
@@ 0.3. Gromov-Witten theory of families of local elliptic curves
Let $\overline{\mathcal{M}}_{1,r}$ be the moduli space of Deligne-Mumford stable curves of genus $1$ with $r$ markings. [^8] We will always assume $r>0$ for stability. Let
@@
be the universal local curve over $\overline{\mathcal{M}}_{1,r}$. The torus $\mathsf{T}=(\mathbb{C}^{*})^{2}$ acts on the $\mathbb{C}^{2}$ factor as before.
Let $\mu^{1},\ldots,\mu^{r}\in\mathrm{Part}(n)$, and let $\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$ be the moduli space of stable [^9] The superscript $\bullet$ indicates possibly disconnected domain curves (but no connected component of the domain is contracted to a point). We follow the conventions of [4]. relative maps to the fibers of $\pi_{\mathbb{C}^{2}}$,
@@
is the moduli space of stable maps of genus $g$ to $E\times\mathbb{C}^{2}$ relative to the divisors determined by the nodes and the markings of $E$ with boundary [^10] The boundary conditions are unordered, and the cohomology weights of the boundary conditions are all the identity class. condition $\mu^{i}$ over the divisor $p_{i}\times\mathbb{C}^{2}$. Since the degree $n$ is recorded in the size of the partitions $\mu^{i}$, we omit $n$ from the notation for $\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$.
The moduli space $\overline{\mathcal{M}}^{\bullet}_{g}(\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r})$ has $\pi_{\mathbb{C}^{2}}$-relative virtual dimension
$nr+\sum_{i=1}^{r}\ell(\mu^{i})\,.$
@@
The moduli space of stable maps $\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}$ is empty unless $g[b]$ is an integer. The virtual class $\Big{[}\overline{\mathcal{M}}^{\bullet}_{g[b]}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}$ is the $\pi_{\mathbb{C}^{2}}$-relative $\mathsf{T}$-equivariant virtual class of the family of relative stable maps to the fibers of $\pi_{\mathbb{C}^{2}}$. We define
@@
$=$ $\int_{\overline{\mathcal{M}}_{1,r}}\epsilon_{*}\left(\Big{[}\overline{\mathcal{M}}^{\bullet}_{g}\big{(}\pi_{\mathbb{C}^{2}},\mu^{1},\ldots,\mu^{r}\big{)}\Big{]}^{vir_{\pi_{\mathbb{C}^{2}}}}\right).$
@@
after the variable change $q=e^{iu}$.
@@ 0.4. Moduli of abelian varieties
be the universal family of principally polarized abelian varieties. We refer the reader to [2] for the foundations of the study of the moduli of abelian varieties.
A general abelian variety $(X,\theta)$ parameterized by $\mathcal{A}_{g}$ has Picard number $1$. The Noether-Lefschetz locus of $\mathcal{A}_{g}$ parameterizes abelian varieties with Picard number at least 2. The simplest irreducible components of the Noether-Lefschetz locus of $\mathcal{A}_{g}$ are related to the geometry of elliptic curves on abelian varieties. For $g\geq 1$ and $n\geq 1$, let
[…]
Let ${\mathsf{NL}}_{g,n}]\in\mathsf{CH}^{g-1}(\mathcal{A}_{g})$ be the associated cycle class.111111 The $g=1$ case is degenerate: $\mathsf{NL}_{1,n}]=[\mathcal{A}_{1}]$ by definition. All Chow groups are taken with $\mathbb{Q}$-coefficients. The following linear combination of components plays a geometrically important role [^12] As usual, $\sigma_{r}(n)=\sum_{d|n}d^{r}$.
$\widetilde{\mathsf{NL}}_{g,n}]=\sum_{n^{\prime}|n}\sigma_{1}\left(\frac{n}{n^{\prime}}\right)[\mathsf{NL}_{g,n^{\prime}}]\ \in\ \mathsf{CH}^{g-1}(\mathcal{A}_{g})\,$,
@@
or, equivalently [^13] Equation (0.4) is correct also for $g=1$ with the convention $\mathsf{NL}_{1,n}]=[\mathcal{A}_{1}]$.,
@@ 0.5.1. Hilbert schemes of points
[…]
@@
[…]
@@
[…]
@@ 0.5.2. Families of local elliptic curves
Let $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)$ be the moduli space of stable [^14] The superscript $\circ$ indicates connected domain curves. relative maps to the fibers of $\pi$,
@@
is the moduli space of stable maps of 1-pointed connected genus $g$ curves to $E$ of degree $n$ relative to the divisors determined by the nodes of $C$. The moduli space $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)$ has $\pi$-relative virtual dimension 2g12g-1 and total virtual dimension $2g$.
@@ 0.5.3. The homomorphism question
Let $\mathcal{M}_{g}^{\mathsf{ct}}$ be the 3g33g-3 dimensional moduli space of genus $g$ curves of compact type, and let
@@ 0.6. Reduction of multi-point series to 1-point series
By [39, Section 4.2], the quantum powers of $D$ generate the quantum cohomology $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$. More precisely, the set
@@
is a basis of $QH^{*}_{\mathsf{T}}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}))$ as a $\mathbb{Q}(t_{1},t_{2},q)$-vector space. Our first reduction result is that the genus 1 Gromov-Witten theory (0.2) of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ can be effectively reduced to 1-point series in the basis (0.7).
@@ 0.7. Reduction to $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$
The reduction of the genus 1 Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to the basic series $\langle D\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ is more subtle. Since the quantum cohomology of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is semisimple, the Givental-Teleman reconstruction result [16, 49] reduces the full higher genus Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to the genus 0 theory together with the calculation of the $\mathsf{R}$-matrix. The $\mathsf{R}$-matrix was proven to be determined [^17] Since the Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ is equivariant, the associated $\mathsf{CohFT}$ is not conformal. by constant map invariants in [46]. The resulting control of the full Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ was used in [46] to prove both the crepant resolution and the GW/DT correspondences associated to the geometry, but was not sufficient to obtain closed form evaluations of any higher genus series (because of the lack of closed forms for the $\mathsf{R}$-matrix).